OpenStudy (samirahdanyel):

Which represents a quadratic function? f(x) = −8x3 − 16x2 − 4x f (x) = x 2 + 2x − 5 f(x) = + 1 f(x) = 0x2 − 9x + 7

jimthompson5910 (jim_thompson5910):

quadratic functions are of the form ax^2+bx+c where 'a' must be nonzero

jimthompson5910 (jim_thompson5910):

the degree (aka largest exponent) is 2

OpenStudy (samirahdanyel):

okay i think its the first one

jimthompson5910 (jim_thompson5910):

no, that's a cubic function

OpenStudy (kittiwitti1):

\[f(x)=-8x^{3}-16x^{2}-4x\]\[f(x)=x^{2}+2x-5\]\(f(x)=+1\) ??? is this missing something?\[f(x)=0x^{2}-9x+7\]

OpenStudy (kittiwitti1):

The leading term in a quadratic equation is a number followed by \(x^{2}\). And the last one doesn't count because 0 times anything is 0 and cancels the whole thing out.\[f(x)=0x^{2}-9x+7\rightarrow f(x)=\cancel{0x^{2}}-9x+7\]

jimthompson5910 (jim_thompson5910):

Example of a quadratic \[\Large -16x^2+100x+140\] notice how the largest exponent is 2 (degree = 2) ------------------------------ \[\Large {\color{red}{-16}}x^2+{\color{green}{100}}x+{\color{blue}{140}}\] is in the form \[\Large {\color{red}{a}}x^2+{\color{green}{b}}x+{\color{blue}{c}}\] where \[\Large \color{red}{a = -16}\] \[\Large \color{green}{b = 100}\] \[\Large \color{blue}{c = 140}\]

OpenStudy (samirahdanyel):

so it would be on of the fractions ?

jimthompson5910 (jim_thompson5910):

you mean something like choice C? no, because you cannot have a variable in the denominator like that

OpenStudy (samirahdanyel):

no not choice c but choice b

jimthompson5910 (jim_thompson5910):

http://www.mathsisfun.com/algebra/polynomials.html look at the section where they describe what makes a polynomial and what doesn't make a polynomial

OpenStudy (kittiwitti1):

\(\frac{4}{x^{2}}\) is \(not\) a quadratic because that would be \(4\times\frac{1}{x^{2}}\) or \(4\times x^{-2}\)

jimthompson5910 (jim_thompson5910):

yes it's choice b because it's in the form ax^2+bx+c in the case of choice B, we have a = 3/4 b = 2 c = -5

OpenStudy (samirahdanyel):

okay I saw what you were talking about on the website

OpenStudy (samirahdanyel):

thank you guys

OpenStudy (kittiwitti1):

No problem. ☺\[\text{Happy Openstudying!}\]